OpenStudy (anonymous):

Use >, <, or = to complete each statement. 14/25 ___ 56% please help!!!

Parth (parthkohli):

Could you express \(\dfrac{56}{100}\) as a fraction?

OpenStudy (anonymous):

it is equal

hero (hero):

Do what @ParthKohli said, then cross multiply: \[\frac{14}{25} \text{___} \frac{56}{100}\] \[(14 \dot\ 100) \text{___} (25 \dot\ 56)\] Put appropriate symbol afterwards

Parth (parthkohli):

Or, alternatively, you could do this:\[\dfrac{14}{25}\cdot \color{#C00}{\dfrac{4}{4}} = \dfrac{?}{?}\]

hero (hero):

@ParthKohli, I was showing a general method for figuring out if two fractions are <, >, or =

Parth (parthkohli):

Yup, yours is better :)

hero (hero):

For example, if you have: \[\frac{2}{3} \text{___} \frac{5}{7}\] You can cross multiply that to get: \[(2)(7) \text{___} (3)(5)\] \[14 \text{<} 15\] Therefore 2/3 is less than 5/7

OpenStudy (anonymous):

WAIT WHY DID HE SAY IT WAS EQUAL??

hero (hero):

Did you cross multiply? (14)(100) ____(25)(56)

hero (hero):

I was only showing an example of course. The example is just an example to show the process. The actual symbol for your problem is not less than of course.

hero (hero):

Do you understand that I was showing you an example which is not the same as your original problem? Examples don't always have the same answer. The purpose of examples is to demonstrate the process, not give you what the answer is.

OpenStudy (anonymous):

IT IS EQUAL!

hero (hero):

Yes, it is equal because 1400 = 1400

OpenStudy (anonymous):

THANK YOU

hero (hero):

@Viviana!, can you figure out the symbol for this one: \[\frac{7}{8} \text{___} \frac{5}{4}\]